Prove that the set of arithmetic functions form an abelian group under the


mapping Dirichlet multiplication with f1≠0

1
Expert's answer
2021-12-13T16:24:47-0500

Let f and g be arithmetic functions. We define the Dirichlet product of f and g by

"(f*g)(n)=\displaystyle \sum_{d|n}f(d)g(n/d)"

Let S be the set of arithmetic functions f such that "f(1)\u22600"

Since "f * g (1) = f (1) g (1)" , then S is closed under Dirichlet multiplication. Furthermore, simple algebraic manipulation shows * to be both commutative and associative.

Let 

"e(n)=\begin{cases}
 1 &\text{if } n=1 \\
 0 &\text{otherwise } 
\end{cases}"

It is clear that "e * f = f" for all arithmetic functions f, so that e is the identity element. Finally, we must show that, given f ∈ S, there exists an f−1∈ S such that "f * f ^{\u22121} = e" . Given f, we will construct f−1 inductively. First, we need "f *f ^{\u22121} (1) = e (1)" , which occurs if and only if

"f (1) f^{ \u22121} (1) = 1" . Since "f (1) \neq 0" , f-1 is uniquely determined. Now, assume that n > 1 and f−1 has been determined for all k < n. Then we have

"f*f^{-1}(n)=\displaystyle \sum_{d|n}f(d)f^{-1}(n/d)\implies -f(1)*f^{-1}(n)=\displaystyle \sum_{d|n,d>1}f(d)f^{-1}(n/d)"

This uniquely determines f-1(n). Thus, we may uniquely determine an f-1  for all f ∈ S.